OpenStudy (anonymous):

Solve for x. -16 = 8x/5 1)x = 10 2)x = -10 3)x = -25 3/5

OpenStudy (phebe):

WELCOME TO OPENSTUDY!! :D

OpenStudy (rushwr):

Show me your work !

OpenStudy (phebe):

2)x = -10

OpenStudy (rushwr):

@Worior9131

OpenStudy (♪chibiterasu):

-16 = 8x/5 Isolate x. Cancel out the 5 by multiplying by 5, since it is being divided. -16 * 5 = (8x/5) * 5 -80 = 8x Now divide 8 on both sides to cancel out the 8 being multiplied. -80/8 = 8x/8 -10 = x

OpenStudy (phebe):

^^^^^^

OpenStudy (anonymous):

what about Solve for x. ⅞ x = -5 (fraction is negative) 1)x = -4 3/8 2)x = 4 3/8 3)x = 5 5/7

OpenStudy (rushwr):

\[-16 = \frac{ 8x }{5 }\] \[-16* 5 = 8x\] \[-80 = 8x\] \[x = \frac{ -80 }{ 8} = -10\]

OpenStudy (rushwr):

Is that the only answers given ?

OpenStudy (rushwr):

And @Worior9131 Can u open up a new question for this ?